Solve the given equation and determine whether the equation is an identity, a conditional equation or an inconsistent equation.

1a. 2(x+5)=4x-2(x-5)

1b. (x/2)-(3/4)=(10-15x)/(6)

1c. 5x-3=5+5x
1
Expert's answer
2013-01-22T09:21:19-0500
1a. 2(x+5)=4x-2(x-5) => 2x+10=4x-2x+10 => 4x+10=4x+10=> 0x=0 => identity equation;
1b. (x/2)-(3/4)=(10-15x)/6 => (2x-3)/4=(10-15x)/6 => 6(2x-3)=4(10-15x)=> 12x-18=40-60x => 72x=58 => x=58/72 => conditional equation;
1c. 5x-3=5+5x => 0x=8 => inconsistent equation;